# Omni2Sound: Towards Unified Video-Text-to-Audio Generation

### v1

@@ 1 Introduction
Early audio generation models typically rely on unimodal conditioning. Text-to-Audio (T2A) [kreuk2022audiogen, liu2023audioldm, evans2024stableaudioopen, ghosal2023instruction] offers strong semantic fidelity and generalization but lacks dense temporal control. Conversely, Video-to-Audio (V2A) [luo2023diff-foley, wang2024frieren, zhang2024foleycraft, chen2024videoguided] ensures fine-grained temporal synchronization with video, yet suffers from weak reasoning in complex scenes and unfaithful generation (e.g., unexpected music or speech) [bain2025vintage, yu2024reasonaudio]. To address this, recent Video-Text-to-Audio (VT2A) methods [chen2024hunyuanfoley, liu2025thinksound, cheng2025mmaudio, tian2025audiox, bain2025vintage] jointly condition on video and text. While VT2A achieves strong both semantic understanding and temporal alignment, its reliance on simultaneous inputs constrains its applicability [tian2025audiox]. Crucially, most VT2A systems lack robustness [liu2025thinksound, chen2024hunyuanfoley, bain2025vintage], degrading sharply under missing-modality conditions (video-only or text-only).
These constraints motivate a unified framework natively supporting VT2A, V2A, and T2A within a single model. This unified paradigm aligns with the AIGC shift, eliminating the redundant architectures and deployment complexity of hard-switching between specialized models. Recent work has begun to advance this unified approach. MMAudio [cheng2025mmaudio] introduces a multimodal joint training framework to improve V2A generation, optionally conditioning on text using large-scale text–audio pairs. Moreover, AudioX [tian2025audiox] enhanced flexibility by supporting broader modality combinations. Despite this progress, two challenges in the unified VT2A framework remain underexplored.  [^1] https://swapforward.github.io/Omni2Sound
@@
First, there is a scarcity of high-quality audio captions that are well-aligned with both audio and video cues. Most unified or specialized VT2A studies create their (V, T, A) training triplets by pairing videos (V) and their audio (A) with captions (T) generated solely from the audio [tian2025audiox, chen2024hunyuanfoley]. However, this approach introduces severe semantic conflict in the multimodal training data (see Figure 1): a frequent mismatch between the visual content and the (audio-only) text caption. This conflict is rooted in the audio modality’s inherent ambiguity (e.g., a tennis hit vs. distant fireworks, or car engine noise vs. an electric drill). This fundamental ambiguity is then exacerbated by the limited capabilities of earlier audio-language models, which are prone to severe hallucinations (e.g., omissions and mislabels) [chen2025detecting]. In our preliminary experiments, we found these modality conflicts caused by mismatches between V-T conditions directly lead to unstable convergence and a significant degradation in audio faithfulness. Unfortunately, there is still a lack of high-quality V-T-A triples for unified VT2A models training, as we further discuss in Section 2.
@@
Second, two critical types of task competition within unified VT2A frameworks remain underexplored. (1) Cross-Task Competition. Prior work, notably MMAudio [cheng2025mmaudio], established that incorporating T-A pairs enhances the generalization and quality of V2A generation. However, training a unified model to excel at both V2A and T2A presents a significant challenge: as shown in our preliminary experiment (Table 5), this joint training introduces a severe T2A-V2A adverse trade-off, rooted in the heterogeneity between text and video modalities. Prioritizing one task during training consistently degrades the performance of the other, indicating a zero-sum optimization dynamic. (2) Intra-Task Competition. We also observe competition within the VT2A task itself. This competition manifests as a modality bias during generation process that undermines cross-conditional consistency, revealing two key failure modes: a bias towards text leads to poor A-V synchronization (Table 6), while a bias towards video exhibits low text-audio faithfulness in off-screen generation scenarios (Table 7).
@@
To address data scarcity, we first introduce SoundAtlas in Section 3, a large-scale, agent-generated multimodal audio-caption dataset. It augments the two largest audio datasets, VGGSound [chen2020vggsound] and AudioSet [gemmeke2017audio], providing semantically rich and temporally detailed captions that even surpass human-expert quality (Table 2). Built on current advanced multimodal foundation models (Gemini-2.5 Pro [team2023gemini] and Qwen-2.5-VL [qwen2025vl]), we develop a multi-turn, agentic annotation pipeline featuring a junior–senior agent handoff, vision-to-language compression, and post-hoc hallucination filtering. This pipeline delivers cost-controlled annotations while maintaining tight visual–audio–text (V–A–T) alignment and a markedly higher text-audio faithfulness than prior datasets. Interestingly, we find its quality is high enough to even correct human annotation errors in VGGSounder [zverev2025vggsounder].
@@
Finally, we construct VGGSound-Omni in Section 5, the first comprehensive benchmark to establish a unified evaluation standard for VT2A, V2A, and T2A. It provides high-quality, human-verified annotations for all three tasks and introduces a challenging off-screen audio generation track. As a result, with a vanilla DiT [Peebles2023DiT] backbone, Omni2Sound achieves unified state-of-the-art performance across all three tasks against both unified and specialized models, showing high-fidelity audio quality, tight audio-visual synchronization, and excellent generation faithfulness.
@@ 2 Related Works
Human-annotated benchmarks like AudioCaps [kim2019audiocaps] (46k) and Clotho [drossos2020clotho] (5k) offer high-quality alignment, but their limited scale, high cost, and lack of detail make them unsuitable for training modern, large-scale models. Automated pipelines emerged to address data scarcity. WavCaps [mei2023wavcaps] used LLMs to refine noisy web metadata (400k captions), and AudioSetCaps [lipping2024audiocaps] used ALMs+LLMs to extract and aggregate details from audio, speech, and music, significantly increasing data volume. As detailed in our Introduction, these audio-only methods suffer from high hallucination rates and can lead to cross-modal conflicts that destabilize VT2A training, stemming from the audio modality’s inherent ambiguity. Visually-enhanced (VE) annotation pipelines like Auto-ACD [han2024autoacd] and Sound-VECaps [yu2024soundvecaps] emerged to leverage visual cues for cross-modal constraint. While promising, existing implementations adopt a separate-then-fuse pipeline: unimodal models extract separate textual cues (e.g., image captions, audio tags), which are then merged by a final LLM. This pipeline is suboptimal, as the LLM fuses lossy textual representations, not raw modalities, leading to the accumulation and amplification of unimodal hallucinations. While using native end-to-end multimodal models (e.g., Gemini [team2023gemini]) seems a natural solution, it also proves suboptimal. As we demonstrate in Section 3, this method faces prohibitive costs and a pervasive visual bias that prevents truly audio-centric captioning. There remains a lack of a large-scale, high-quality visual–audio–text (V–A–T) aligned audio caption dataset suitable for training unified VT2A models.
@@
The audio generation paradigm is shifting towards unified, omni-modal frameworks, a trajectory initiated by MMAudio [cheng2025mmaudio]. While it integrated V2A and T2A, its approach was fundamentally V2A-centric, using T-A pairs merely as augmentation for V2A rather than optimizing T2A as a co-equal task. Subsequent works like AudioX [tian2025audiox] and AudioGen-Omni [wang2025audiogenomni] expanded this scope to more flexible modality combinations. However, these efforts often relied on brute-force data scaling (e.g., AudioX with over 9 million samples), which revealed inefficiencies and failed to yield proportional SOTA performance. Critically, these early models [cheng2025mmaudio, tian2025audiox, wang2025audiogenomni] largely overlooked the inherent cross-task competition stemming from co-training these diverse sub-tasks. UniFlow-Audio [xu2025uniflow] is the first to systematically address this by categorizing tasks into Time-Aligned (TA) and Non-Time-Aligned (NTA) classes and analyzing their competitive dynamics. However, its analysis remains coarse-grained, failing to investigate the granular competition within the TA category (i.e., V2A vs. T2A). Moreover, the challenging case of joint cross-modal generation (VT2A) remains unaddressed. Consequently, a fundamental study on task competitive dynamics within a unified VT2A framework remains absent.
@@ 3 SoundAtlas: V-A-T Data Construction
Existing automated audio caption datasets often suffer from severe visual-audio-text (V-A-T) misalignment with high hallucination rates due to the limitations of early ALMs [lipping2024audiocaps, han2024autoacd, yu2024soundvecaps]. While recent native multimodal foundation models like Gemini 2.5 [team2023gemini, MMAR:2025, Qwen3Omni:2025] offer strong capabilities, we find that a naive implementation—processing raw video-audio pairs directly—is suboptimal for audio caption dataset construction. Specifically, it incurs prohibitive costs (approx. \$10,275 per 1M samples; see Appendix A) and suffers from inherent visual bias, where models hallucinate auditory labels for non-existent events due to visual interference, as shown in Figure 1.
@@
We first apply A-V Consistency Routing on raw video from AudioSet [gemmeke2017audio] and VGGSound [chen2020vggsound]. This step is based on the core finding that visual cues are reliable for high-consistency A-V clips but act as distractors in low-consistency clips as shown in Figure 2. We classify samples based on ImageBind alignment ($s_{ib}$) using thresholds $\tau_{\text{low}}=0.20$ and $\tau_{\text{high}}=0.30$: (i) High-consistency ($s_{ib}>\tau_{\text{high}}$) enter the A-V Enhanced Path; (ii) Medium-consistency ($\tau_{\text{low}}\leq s_{ib}\leq\tau_{\text{high}}$) are routed to the Audio-Only Path to prevent visual hallucinations; and (iii) Noise ($s_{ib}<\tau_{\text{low}}$) is discarded.
@@
AudioSetCaps [lipping2024audiocaps] 0.330 0.397 0.351 0.421
Sound-VECaps [yu2024soundvecaps] 0.370 0.425 - -
Auto-ACD [han2024autoacd] 0.396 0.437 0.409 0.457
@@
Table 2: Caption quality comparison via MLLM-as-a-judge and human evaluation, reporting the Mean Win Rate for Semantic (MWR-S) and Temporal (MWR-T) alignment. Human-Expert refers to the human-annotated captions from AudioCaps [kim2019audiocaps].
@@
Auto-ACD [han2024autoacd] 0.39 0.41 0.31 0.26
Human-Expert [kim2019audiocaps] 0.36 0.51 0.46 0.55
@@
This step implements our key insight: vision must be treated as a contextual constraint, not a primary input. We found that compressing the visual stream into a textual representation ($c_{v}$) is a more effective strategy, as it simultaneously addresses both of our defined challenges. First, it addresses cost by replacing the prohibitively expensive raw video input ($V+A$) with a cost-effective text-audio prompt ($c_{v}+A$). Second, it robustly mitigates cross-modal hallucinations by filtering the visual bias, providing only low-bias semantic context (e.g., "A man and a woman are standing…") rather than a misleading raw visual stream. Therefore, for samples $V$ routed to the A-V Enhanced Path, we use Qwen-2.5-VL [qwen2025vl] to analyze the video $V$ (without its audio $A$) and generate the textual representation $c_{v}=\text{Qwen}(V)$. Conversely, samples on the Audio-Only Path are assigned a null context.
@@
All samples then enter our handoff pipeline. The task is first assigned to the Junior agent, $G_{\text{junior}}$ (Gemini 2.5 Flash), which receives the audio $A$ and the optional visual context $c_{v}$. Let the output caption be $c_{a}=G_{\text{junior}}(A,c_{v})$. This caption $c_{a}$ is then flagged if it (i) meets our complexity criteria (text-based heuristics to identify complex audio scenes), (ii) contains high-frequency hallucination phrases, or (iii) fails our differentiated CLAP [elizalde2023clap] check, $\text{CLAP}(c_{a},A)<\tau_{clap}$, where $\tau_{clap}$ is $0.35$ for general audio and $0.15$ for music. Flagged tasks are escalated to the Senior agent, $G_{\text{senior}}$ (Gemini 2.5 Pro). To control costs, the Senior agent’s reasoning output is limited to 128 tokens, providing a more precise caption.
@@
Finally, all generated captions $c_{a}$ undergo a two-stage verification. First, a CLAP (T-A) filtering model [elizalde2023clap] ensures high Text-Audio faithfulness; captions where $\text{CLAP}(c_{a},A)<\tau_{verify}$ are discarded. Second, for captions from the A-V Enhanced Path ($c_{v}\neq\emptyset$), an A-V-T Verifier, $V_{\text{AVT}}$, ensures $c_{a}$ is a reasonable acoustic inference given $c_{v}$. Captions that pass all filters are accepted into the final dataset $\mathcal{D}_{\text{SoundAtlas (Ours)}}$, which augments VGGSound [chen2020vggsound] and AudioSet [gemmeke2017audio] datasets with human-expert-level audio captions.
@@ 3.1 Comparison with Existing Pipeline
We compare SoundAtlas against other automated pipelines [lipping2024audiocaps, han2024autoacd, yu2024soundvecaps] on high audio-visual consistency subsets sourced from AudioSet and VGGSound, where ImageBind score $s_{ib}>0.30$. As shown in Table 1, SoundAtlas significantly outperforms all competitors on both LA-CLAP and MS-CLAP scores, demonstrating superior text-audio alignment. Additionally, we conduct a fine-grained MLLM-as-a-judge (Gemini 3.0 Pro [team2023gemini]) evaluation on the intersection of AudioCaps and all compared datasets [kim2019audiocaps]. As shown in Table 2, SoundAtlas achieves a substantially higher mean win rate in semantic alignment (MWR-S) and temporal alignment (MWR-T) than both the strongest baseline (Auto-ACD) and the Human-Expert annotations, across both semantic and temporal alignment. To mitigate potential evaluation bias, a follow-up human validation study is conducted, further corroborating our results (details in Appendix Section C). As illustrated in Figure 2 (right), SoundAtlas demonstrates clear superiority over existing automated datasets, characterized by its richer semantic content and explicit temporal ordering.

### v2

@@ 1 Introduction
Early audio generation models typically rely on unimodal conditioning. Text-to-Audio (T2A) [27, 30, 13, 16] offers strong semantic fidelity and generalization but lacks dense temporal control. Conversely, Video-to-Audio (V2A) [33, 47, 53, 6] ensures fine-grained temporal synchronization with video, yet suffers from weak reasoning in complex scenes and unfaithful generation (e.g., unexpected music or speech) [29, 21]. To address this, recent Video-Text-to-Audio (VT2A) methods [40, 32, 7, 43, 29] jointly condition on video and text. While VT2A achieves strong both semantic understanding and temporal alignment, its reliance on simultaneous inputs constrains its applicability [43]. Crucially, most VT2A systems lack robustness [32, 40, 29], degrading sharply under missing-modality conditions (video-only or text-only).
These constraints motivate a unified framework natively supporting VT2A, V2A, and T2A within a single model. This unified paradigm aligns with the AIGC shift, eliminating the redundant architectures and deployment complexity of hard-switching between specialized models. Recent work has begun to advance this unified approach. MMAudio [7] introduces a multimodal joint training framework to improve V2A generation, optionally conditioning on text using large-scale text–audio pairs. Moreover, AudioX [43] enhanced flexibility by supporting broader modality combinations. Despite this progress, two challenges in the unified VT2A framework remain underexplored.  [^1] https://swapforward.github.io/Omni2Sound
@@
First, there is a scarcity of high-quality audio captions that are well-aligned with both audio and video cues. Most unified or specialized VT2A studies create their (V, T, A) training triplets by pairing videos (V) and their audio (A) with captions (T) generated solely from the audio [43, 40]. However, this approach introduces severe semantic conflict in the multimodal training data (see Figure 1): a frequent mismatch between the visual content and the (audio-only) text caption. This conflict is rooted in the audio modality’s inherent ambiguity (e.g., a tennis hit vs. distant fireworks, or car engine noise vs. an electric drill). This fundamental ambiguity is then exacerbated by the limited capabilities of earlier audio-language models, which are prone to severe hallucinations (e.g., omissions and mislabels) [5]. In our preliminary experiments, we found these modality conflicts caused by mismatches between V-T conditions directly lead to unstable convergence and a significant degradation in audio faithfulness. Unfortunately, there is still a lack of high-quality V-T-A triples for unified VT2A models training, as we further discuss in Section 2.
@@
Second, two critical types of task competition within unified VT2A frameworks remain underexplored. (1) Cross-Task Competition. Prior work, notably MMAudio [7], established that incorporating T-A pairs enhances the generalization and quality of V2A generation. However, training a unified model to excel at both V2A and T2A presents a significant challenge: as shown in our preliminary experiment (Table 5), this joint training introduces a severe T2A-V2A adverse trade-off, rooted in the heterogeneity between text and video modalities. Prioritizing one task during training consistently degrades the performance of the other, indicating a zero-sum optimization dynamic. (2) Intra-Task Competition. We also observe competition within the VT2A task itself. This competition manifests as a modality bias during generation process that undermines cross-conditional consistency, revealing two key failure modes: a bias towards text leads to poor A-V synchronization (Table 6), while a bias towards video exhibits low text-audio faithfulness in off-screen generation scenarios (Table 7).
@@
To address data scarcity, we first introduce SoundAtlas in Section 3, a large-scale, agent-generated multimodal audio-caption dataset. It augments the two largest audio datasets, VGGSound [4] and AudioSet [15], providing semantically rich and temporally detailed captions that even surpass human-expert quality (Table 2). Built on current advanced multimodal foundation models (Gemini-2.5 Pro [42] and Qwen-2.5-VL [28]), we develop a multi-turn, agentic annotation pipeline featuring a junior–senior agent handoff, vision-to-language compression, and post-hoc hallucination filtering. This pipeline delivers cost-controlled annotations while maintaining tight visual–audio–text (V–A–T) alignment and a markedly higher text-audio faithfulness than prior datasets. Interestingly, we find its quality is high enough to even correct human annotation errors in VGGSounder [54].
@@
Finally, we construct VGGSound-Omni in Section 5, the first comprehensive benchmark to establish a unified evaluation standard for VT2A, V2A, and T2A. It provides high-quality, human-verified annotations for all three tasks and introduces a challenging off-screen audio generation track. As a result, with a vanilla DiT [37] backbone, Omni2Sound achieves unified state-of-the-art performance across all three tasks against both unified and specialized models, showing high-fidelity audio quality, tight audio-visual synchronization, and excellent generation faithfulness.
@@ 2 Related Works
Human-annotated benchmarks like AudioCaps [24] (46k) and Clotho [10] (5k) offer high-quality alignment, but their limited scale, high cost, and lack of detail make them unsuitable for training modern, large-scale models. Automated pipelines emerged to address data scarcity. WavCaps [36] used LLMs to refine noisy web metadata (400k captions), and AudioSetCaps [2] used ALMs+LLMs to extract and aggregate details from audio, speech, and music, significantly increasing data volume. As detailed in our Introduction, these audio-only methods suffer from high hallucination rates and can lead to cross-modal conflicts that destabilize VT2A training, stemming from the audio modality’s inherent ambiguity. Visually-enhanced (VE) annotation pipelines like Auto-ACD [41] and Sound-VECaps [51] emerged to leverage visual cues for cross-modal constraint. While promising, existing implementations adopt a separate-then-fuse pipeline: unimodal models extract separate textual cues (e.g., image captions, audio tags), which are then merged by a final LLM. This pipeline is suboptimal, as the LLM fuses lossy textual representations, not raw modalities, leading to the accumulation and amplification of unimodal hallucinations. While using native end-to-end multimodal models (e.g., Gemini [42]) seems a natural solution, it also proves suboptimal. As we demonstrate in Section 3, this method faces prohibitive costs and a pervasive visual bias that prevents truly audio-centric captioning. There remains a lack of a large-scale, high-quality visual–audio–text (V–A–T) aligned audio caption dataset suitable for training unified VT2A models.
@@
The audio generation paradigm is shifting towards unified, omni-modal frameworks, a trajectory initiated by MMAudio [7]. While it integrated V2A and T2A, its approach was fundamentally V2A-centric, using T-A pairs merely as augmentation for V2A rather than optimizing T2A as a co-equal task. Subsequent works like AudioX [43] and AudioGen-Omni [46] expanded this scope to more flexible modality combinations. However, these efforts often relied on brute-force data scaling (e.g., AudioX with over 9 million samples), which revealed inefficiencies and failed to yield proportional SOTA performance. Critically, these early models [7, 43, 46] largely overlooked the inherent cross-task competition stemming from co-training these diverse sub-tasks. UniFlow-Audio [50] is the first to systematically address this by categorizing tasks into Time-Aligned (TA) and Non-Time-Aligned (NTA) classes and analyzing their competitive dynamics. However, its analysis remains coarse-grained, failing to investigate the granular competition within the TA category (i.e., V2A vs. T2A). Moreover, the challenging case of joint cross-modal generation (VT2A) remains unaddressed. Consequently, a fundamental study on task competitive dynamics within a unified VT2A framework remains absent.
@@ 3 SoundAtlas: V-A-T Data Construction
Existing automated audio caption datasets often suffer from severe visual-audio-text (V-A-T) misalignment with high hallucination rates due to the limitations of early ALMs [2, 41, 51]. While recent native multimodal foundation models like Gemini 2.5 [42, 34, 49] offer strong capabilities, we find that a naive implementation—processing raw video-audio pairs directly—is suboptimal for audio caption dataset construction. Specifically, it incurs prohibitive costs (approx. \$10,275 per 1M samples; see Appendix A) and suffers from inherent visual bias, where models hallucinate auditory labels for non-existent events due to visual interference, as shown in Figure 1.
@@
We first apply A-V Consistency Routing on raw video from AudioSet [15] and VGGSound [4]. This step is based on the core finding that visual cues are reliable for high-consistency A-V clips but act as distractors in low-consistency clips as shown in Figure 2. We classify samples based on ImageBind alignment ($s_{ib}$) using thresholds $\tau_{\text{low}}=0.20$ and $\tau_{\text{high}}=0.30$: (i) High-consistency ($s_{ib}>\tau_{\text{high}}$) enter the A-V Enhanced Path; (ii) Medium-consistency ($\tau_{\text{low}}\leq s_{ib}\leq\tau_{\text{high}}$) are routed to the Audio-Only Path to prevent visual hallucinations; and (iii) Noise ($s_{ib}<\tau_{\text{low}}$) is discarded.
@@
AudioSetCaps [2] 0.330 0.397 0.351 0.421
Sound-VECaps [51] 0.370 0.425 - -
Auto-ACD [41] 0.396 0.437 0.409 0.457
@@
Table 2: Caption quality comparison via MLLM-as-a-judge and human evaluation, reporting the Mean Win Rate for Semantic (MWR-S) and Temporal (MWR-T) alignment. Human-Expert refers to the human-annotated captions from AudioCaps [24].
@@
Auto-ACD [41] 0.39 0.41 0.31 0.26
Human-Expert [24] 0.36 0.51 0.46 0.55
@@
This step implements our key insight: vision must be treated as a contextual constraint, not a primary input. We found that compressing the visual stream into a textual representation ($c_{v}$) is a more effective strategy, as it simultaneously addresses both of our defined challenges. First, it addresses cost by replacing the prohibitively expensive raw video input ($V+A$) with a cost-effective text-audio prompt ($c_{v}+A$). Second, it robustly mitigates cross-modal hallucinations by filtering the visual bias, providing only low-bias semantic context (e.g., "A man and a woman are standing…") rather than a misleading raw visual stream. Therefore, for samples $V$ routed to the A-V Enhanced Path, we use Qwen-2.5-VL [28] to analyze the video $V$ (without its audio $A$) and generate the textual representation $c_{v}=\text{Qwen}(V)$. Conversely, samples on the Audio-Only Path are assigned a null context.
@@
All samples then enter our handoff pipeline. The task is first assigned to the Junior agent, $G_{\text{junior}}$ (Gemini 2.5 Flash), which receives the audio $A$ and the optional visual context $c_{v}$. Let the output caption be $c_{a}=G_{\text{junior}}(A,c_{v})$. This caption $c_{a}$ is then flagged if it (i) meets our complexity criteria (text-based heuristics to identify complex audio scenes), (ii) contains high-frequency hallucination phrases, or (iii) fails our differentiated CLAP [11] check, $\text{CLAP}(c_{a},A)<\tau_{clap}$, where $\tau_{clap}$ is $0.35$ for general audio and $0.15$ for music. Flagged tasks are escalated to the Senior agent, $G_{\text{senior}}$ (Gemini 2.5 Pro). To control costs, the Senior agent’s reasoning output is limited to 128 tokens, providing a more precise caption.
@@
Finally, all generated captions $c_{a}$ undergo a two-stage verification. First, a CLAP (T-A) filtering model [11] ensures high Text-Audio faithfulness; captions where $\text{CLAP}(c_{a},A)<\tau_{verify}$ are discarded. Second, for captions from the A-V Enhanced Path ($c_{v}\neq\emptyset$), an A-V-T Verifier, $V_{\text{AVT}}$, ensures $c_{a}$ is a reasonable acoustic inference given $c_{v}$. Captions that pass all filters are accepted into the final dataset $\mathcal{D}_{\text{SoundAtlas (Ours)}}$, which augments VGGSound [4] and AudioSet [15] datasets with human-expert-level audio captions.
@@ 3.1 Comparison with Existing Pipeline
We compare SoundAtlas against other automated pipelines [2, 41, 51] on high audio-visual consistency subsets sourced from AudioSet and VGGSound, where ImageBind score $s_{ib}>0.30$. As shown in Table 1, SoundAtlas significantly outperforms all competitors on both LA-CLAP and MS-CLAP scores, demonstrating superior text-audio alignment. Additionally, we conduct a fine-grained MLLM-as-a-judge (Gemini 3.0 Pro [42]) evaluation on the intersection of AudioCaps and all compared datasets [24]. As shown in Table 2, SoundAtlas achieves a substantially higher mean win rate in semantic alignment (MWR-S) and temporal alignment (MWR-T) than both the strongest baseline (Auto-ACD) and the Human-Expert annotations, across both semantic and temporal alignment. To mitigate potential evaluation bias, a follow-up human validation study is conducted, further corroborating our results (details in Appendix Section C). As illustrated in Figure 2 (right), SoundAtlas demonstrates clear superiority over existing automated datasets, characterized by its richer semantic content and explicit temporal ordering.